OpenStudy (anonymous):

m<3=2y m<4= x+y m<5= 2x-y <3 is congruent to <5

OpenStudy (anonymous):

is this yur question m♥=2y

jimthompson5910 (jim_thompson5910):

angle 3 is congruent to angle 5, so angle 3 = angle 5 2y = 2x - y 2y + y = 2x 3y = 2x

OpenStudy (anonymous):

i need x and y simplifyed though

jimthompson5910 (jim_thompson5910):

does it mention how angle 4 is connected at all?

jimthompson5910 (jim_thompson5910):

or is there a pic showing how angles 3, 4 and 5 are connected?

OpenStudy (anonymous):

3 is supplamentary to 4

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

oh i picked the wrong angle

jimthompson5910 (jim_thompson5910):

3 is supplementary to 4 angle 3 + angle 4 = 180 2y + x + y = 180 x + 3y = 180

jimthompson5910 (jim_thompson5910):

earlier we found that 3y = 2x, so plug this in to get x + 3y = 180 x + 2x = 180 ... plug in 3y = 2x 3x = 180 x = 180/3 x = 60

jimthompson5910 (jim_thompson5910):

then use this to find y 3y = 2x 3y = 2*60 3y = 120 y = 120/3 y = 40

OpenStudy (anonymous):

THANK YOU

jimthompson5910 (jim_thompson5910):

you're welcome